OpenStudy (anonymous):

Find the first four terms of the sequence represented by the expression. 1. 5+2n 2. 2-2n 3. -7n 4. 5n+6

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

well i dont know how to calculate it

Mehek (mehek14):

ok so to get the first term, plug in 1 for n for second term plug in 2 for n for third term plug in 3 for n for fourth term plug in 4 for n

Mehek (mehek14):

so for #1 5+2n 5 + 2*1 = ?

OpenStudy (anonymous):

7

Mehek (mehek14):

that's the first term

Mehek (mehek14):

now for the 2nd term 5 + 2*2 = 3rd 5 + 2*3 = 4th 5 + 2*4 =

OpenStudy (anonymous):

2nd=9 3rd=11 4th=13 am i right @Mehek14

Mehek (mehek14):

yes so that would be for number 1

Mehek (mehek14):

7 , 9 , 11 , 13

Mehek (mehek14):

will you be able to do the rest on your own?

OpenStudy (anonymous):

i can try thank you here is your medal

Mehek (mehek14):

thanks and gl :)

OpenStudy (anonymous):

i cant do number 2

Mehek (mehek14):

if you want just do them all and i can check

Mehek (mehek14):

2-2n 1st 2 - 2*1 = 2nd 2 - 2*2 = 3rd 2 - 2*3 = 4th 2 - 2*4 =

OpenStudy (anonymous):

1st=0 2nd=-2 3rd=-4 4th=-6

Mehek (mehek14):

correct

OpenStudy (anonymous):

1st=-7 2nd=-14 3rd=-21 4th=-28

Mehek (mehek14):

good

OpenStudy (anonymous):

1st=11 2nd=16 3rd=21 4th=26

Mehek (mehek14):

good job \(\ddot \smile\)

OpenStudy (anonymous):

i got ten more questions to go wanna help me

Mehek (mehek14):

can you post separately and idk if i have time for all of them

OpenStudy (anonymous):

i can but you can still help me with a couple of them

Mehek (mehek14):

ok